OpenStudy (anonymous):

Simplify by using the Distributive Property and combining of similar terms: 5(x + 3) - 4(2x - 1)

OpenStudy (whpalmer4):

Distributive property: \[a(b+c) = ab + ac\] Use that to expand the expression, then combine similar terms.

Directrix (directrix):

5(x + 3) - 4(2x - 1) = 5*x + 5*3 - 4*(2x) -4 * (-1) What is the next step? @shell1903

OpenStudy (anonymous):

I came up with -3x+19

Directrix (directrix):

@shell1903 Correct.